I am taking a practice test. One of my questions is: "A rancher is planning to put up 220 yards of fencing. In the morning, she puts up 80 yards of fencing. In the afternoon she puts up 40% of the remaining fence. What is the total percentage of the amount of fencing she has put up today?"
I know the answer is 62%. Every time i work out the answer however, i get 76%. How do you work this problem out?

Found 2 solutions by Alan3354, rfer:
Answer by Alan3354(69443)

"A rancher is planning to put up 220 yards of fencing. In the morning, she puts up 80 yards of fencing. In the afternoon she puts up 40% of the remaining fence. What is the total percentage of the amount of fencing she has put up today?"
I know the answer is 62%. Every time i work out the answer however, i get 76%. How do you work this problem out?
-------------------------
80 yds in the morning leaves 140 yds
40% of 140 = 56 yds, for a total of 136 yds
136/220 = 0.618 =~ 62%

Answer by rfer(16322)

220-80=140
140*.4=56
80+56=136
136/220=0.618or rounded to 62%